# references.bib AIJIM: A Theoretical Model for Real-Time, Crowdsourced Environmental Journalism with AI

### v3

@@ VI Conclusion and Future Work
The Artificial Intelligence Journalism Integration Model (AIJIM) redefines environmental journalism by seamlessly integrating real-time hazard detection, AI-driven reporting (e.g., via GPT-4 in this study), and crowdsourced validation. The 2024 Mallorca pilot on the NamicGreen platform showcased its prowess: achieving an 85.4% detection accuracy with a DETR-based model, uncovering 50 previously undocumented illegal waste sites, slashing reporting latency by 40% using GPT-4-generated reports, and mobilizing 252 citizen validators [Tiltack2025]. These results, achieved with one possible configuration, mark a leap forward in speed, precision, and engagement over traditional methods (Table III), positioning AIJIM as a pioneering, adaptable force in AI-driven journalism.
@@
VI-A Practical and Scientific Implications
@@
VI-B Future Development Priorities
@@
VI-C AI, Journalism, and Societal Impact
@@
VI-D International Regulatory Considerations
@@
VI-E Conclusion
AIJIM exemplifies AI’s revolutionary potential in environmental journalism, delivering a real-time, scalable, and ethical solution for global hazard detection. By explicitly integrating dual-level human validation—crowdsourced visual hazard verification and human review of AI-generated textual reports—with AI-driven insights, AIJIM ensures comprehensive accuracy and transparency. Demonstrated through the Mallorca pilot configuration, AIJIM does not just enhance journalism—it redefines its very foundation, shaping a future where technology and citizen collaboration drive impactful, evidence-based environmental action adaptable to various models and contexts [Tiltack2025].
Appendix A Detailed Tables
@@
| Bias in Rural Detection Accuracy | 30% lower accuracy due to dataset limitations |
@@
| AI Accuracy | Precision (85.4%) outpaces manual methods [Tiltack2025] | Rural bias (30% drop) requires dataset diversity |
| Trust | XAI (SHAP, LIME) boosts transparency [Shin2021] | Black-box concerns demand clearer accountability |
| Engagement | 252 validators enhance reliability and inclusion [Verma2024] | Misinformation risks need robust controls |
| Regulation | GDPR and EU AI Act alignment [Gstrein2024] | Accountability for AI outputs remains unresolved |
@@
Appendix B Appendix: Visual Examples of AIJIM Implementation

### v5

@@ 6 Conclusion and Future Work
The Artificial Intelligence Journalism Integration Model (AIJIM) redefines environmental journalism by seamlessly integrating real-time hazard detection, AI-driven reporting (e.g., via GPT-4 in this study), and crowdsourced validation. The 2024 Mallorca pilot on the NamicGreen platform showcased its prowess: achieving an 85.4% detection accuracy with a DETR-based model, uncovering 50 previously undocumented illegal waste sites, slashing reporting latency by 40% using GPT-4-generated reports, and mobilizing 252 citizen validators [34]. These results, achieved with one possible configuration, mark a leap forward in speed, precision, and engagement over traditional methods (Table 3), positioning AIJIM as a pioneering, adaptable force in AI-driven journalism.
@@
6.1 Practical and Scientific Implications
@@
6.2 Future Development Priorities
@@
6.3 AI, Journalism, and Societal Impact
@@
6.4 International Regulatory Considerations
@@
6.5 Conclusion
AIJIM exemplifies AI’s revolutionary potential in environmental journalism, delivering a real-time, scalable, and ethical solution for global hazard detection. By explicitly integrating dual-level human validation—crowdsourced visual hazard verification and human review of AI-generated textual reports—with AI-driven insights, AIJIM ensures comprehensive accuracy and transparency. Demonstrated through the Mallorca pilot configuration, AIJIM does not just enhance journalism—it redefines its very foundation, shaping a future where technology and citizen collaboration drive impactful, evidence-based environmental action adaptable to various models and contexts [34].
Explainability Trade-offs and Practical Implications
AIJIM’s explainability strategy combines fast visual attribution through CAM overlays with the option of model-agnostic, per-box explanations using LIME. This dual-layer approach provides both efficiency for real-time applications and depth for expert-level validation. In practice, the CAM method supports immediate visual feedback with low computational cost, while LIME explanations—available via an additional API—enable in-depth audits and detailed traceability when needed. This balance ensures that AIJIM remains scalable without compromising interpretability in high-stakes scenarios.
Explainability Trade-offs and Practical Implications
AIJIM’s explainability architecture balances real-time performance with interpretability through a two-layer approach. The default CAM-based visualization is computationally efficient and integrates directly into the detection pipeline. It has proven useful for general user feedback and field-level reporting.
On the other hand, the LIME module—though slower—is integrated as an optional, asynchronous feature for per-box inspection. It offers localized, model-agnostic insight that can support expert reviews and auditing processes. Together, these layers provide a practical trade-off between speed and explanatory depth, aligning well with AIJIM’s real-world requirements.
Appendix A Appendix A: Detailed Tables
@@
Bias in Rural Detection Accuracy | Slight variation due to dataset and environmental factors, addressed through expanded datasets and adaptive learning [34] |
@@
AI Accuracy | Precision (85.4%) outpaces manual methods [34] | Slight rural performance variation requires dataset diversity [34] |
Trust | XAI (SHAP, LIME) boosts transparency [32] | Black-box concerns demand clearer accountability |
Engagement | 252 validators enhance reliability and inclusion [36] | Misinformation risks need robust controls |
Regulation | GDPR and EU AI Act alignment [19] | Accountability for AI outputs remains unresolved |
@@
Appendix B Appendix B: Visual Examples of AIJIM Implementation